Systems Staff Meeting Notes 4/30/01

Systems Staff Meeting Notes
April 30, 2001

Present: Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Bobbie Meng has returned after a brief interlude at IS in Tigert.
  2. Will is attending W2K training at the Livestock Pavilion for the next four days.
  3. Anguished cry from Personnel – timecards not turned in on time, not signed, incorrectly filled out, etc. As far as we know, Systems is not a culprit, but Bill urges that we continue to keep up the good work.
  4. Winston is a master plumber – he installed a new sink in his garage over the weekend.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders are moving – Web server & Fluke meter have been ordered. Time frame for arrival is at the mercy of Purchasing.
  2. Geoffrey reports an old Gateway 155 in LAC is broken – it is beyond warranty. LAC needs to send in a request for a new machine via regular channels (hardware request form). We do not have spares available.
Liaisons
  1. Bill has a meeting with the PSITC group tomorrow to discuss the future use of Centurion Guard and WinShield with W2K. They have submitted a document to maintain the status quo, yet they want to upgrade the applications.
  2. Administration is requiring that all Library personnel attend two meetings at the Reitz Union this week – one on Wednesday morning, 10 AM, featuring Charles Frazier. The second is on Thursday at 1 PM and David Colburn will be speaking. Each session will last approximately an hour and a half.
Networking
  1. NT11 (the Ref Express server) – NERDC determined that it had been hit by hackers and decided that it had to be nuked. Then they determined that it needed to be rebuilt from scratch. The problem was finally resolved on Friday evening and Ref Express was back up. This was after 36 hours dealing with E-Share product support. An application, CSR.exe, was constantly running – it ate all of the virtual memory and caused the machine to crash - usually caused by connecting with SQL server. There are multiple layers of permissions. Chat, sign-ons, admin and the database use SQL server security, not NT. Net Agent server looks like it should work with NT, but it doesn’t. Bill says it’s some kind of bug that they haven’t addressed yet. There was also an email glitch with the server this morning. It does not use the Exchange server, but NT11. Bill will be meeting with NERDC about this incident. There was no sign of a hacker, no threat – so why did they tell us to rebuild the server?
  2. The electric installation is ongoing in the new server room (tape vault) today. Dell is scheduled to show next Monday, but Will is going to try to move that back a week.
  3. Security system – two ports were needed for a last minute install in AFA (OK), Music & Education. The order for the servers was submitted in March and was on hold in Purchasing for most of April. Supposedly the order has gone out, it does not show on the Dell Order Watch yet – all questions regarding the status should be referred to Business Services.
  4. East 205 closet – phone folks stole our wire from the punch down. There is also wire from the punch down in use by the phone folks in room 129 West. Will contact and have them remove.
  5. Working phones in West closet? Will to check on it.
  6. Room 124 West closet now has air vents. The temperature is currently at 84 degrees. Uncertain if it is just in the process of cooling down or if the vents are too small. Did they pull a permit? Will give it a few days to see if it cools down anymore.
  7. Room 100 – desk damaged in shipment, according to Gary Cornwell. He is dealing with Facilities to have it returned.
Software Projects Status
  1. OPS Tracking changes are done. Phek Su, Lawan and Tim are all happy – they find it easier to use. They will change over to the new version this afternoon.
  2. Personnel database is ongoing.
  3. Elected Official database is complete - all that is left is for Cynthia to install the viewer on the server.
  4. Hilton is working on Library Statistic changes for Jack Waters.
  5. Cynthia is working on PAULA, so that the current desktop version is pointing to the same place as the Web version.
  6. Ringling is finished.
  7. Ying – DLC Tracking database – she has status meetings with Erich every two weeks, but Bill would like to report the modules as finished as it goes, with changes to start a new version.
  8. ILLIAD – starting to bubble for ILL. Order has been placed.
  9. Bobbie mentioned PAULA to Tigert IS and they were very interested. We’ll wait for an official request.
DP Coordinator
  1. Three new students have been hired for the summer semester.
  2. Finals are this week – students will drop off by Friday.
  3. Students have been working on the tape racks. All of the ‘O’ rack-slots are empty, currently about halfway through with ‘N’. Have ‘M’ left to go then the whole first rack can be removed. Electricians will impede progress today. Since the students will be gone next week, Debra will work on it.


Back to the Systems Homepage

Last updated May 7, 2001
by Debra Harris
debharr@mail.uflib.ufl.edu